In Writing explain how to find the asymptote of y=- 3/x-2 +11
1
Expert's answer
2020-04-28T18:47:31-0400

The equations of slant asymptotes are usually sought in the form y = kx + b. By definition, asymptotes:

"\lim\limits_{x\rarr \infty}(kx+b- f(x))\\"

We find the coefficient k:

"k=\lim\limits_{x\rarr \infty}\frac{f(x)}{x}\\
k=\lim\limits_{x\rarr \infty}\frac{\frac{-3}{x-2}+11}{x}=\lim\limits_{x\rarr \infty}\frac{11x-25}{x^2-2x}=0\\"

We find the coefficient b:

"b=\lim\limits_{x\rarr \infty}{f(x)}-{kx}\\
b=\lim\limits_{x\rarr \infty}\frac{-3}{x-2}+11-0x=\lim\limits_{x\rarr \infty}\frac{11x-25}{x-2}=11\\"

We obtain the equation of horizontal asymptote:

"y=11\\"

Find the vertical asymptotes. To do this, define the points of discontinuity:

x1=2

Find the asymptotic behaviour at x = 2:

"\lim\limits_{x\rarr2+0 }\frac{-3}{x-2}+11=-\infty\\
\lim\limits_{x\rarr2-0 }\frac{-3}{x-2}+11=+\infty"

x1 = 2 is a discontinuity point of the second kind and x=2 is a vertical asymptote.